OpenStudy (anonymous):

Determine the domain of the function. f(x) = sqrt(x+1) / (x+4)(x-6)

OpenStudy (anonymous):

\[f(x) = \frac{ \sqrt{x+1} }{ (x+4)(x-6) }\]

OpenStudy (anonymous):

well denominator cant be zero nor you cant have a negative sqr root

OpenStudy (anonymous):

all you have to do is this x+1=>0 x+4=0 x-6=0

OpenStudy (anonymous):

x ≥ -1, x ≠ 6 All real numbers except -4 and 6 All real numbers x ≥ 0

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

well we solve x+1=>0 x>-1 x has to be greater than -1 x+4=0 x=-4 x cant be -4 x-6=0 x=6 x cant be 6

OpenStudy (anonymous):

for the first one x=>-1

OpenStudy (anonymous):

x ≥ -1, x ≠ 6 and x ≠ -4

OpenStudy (anonymous):

I think its A but you are missing the x cant be -4

OpenStudy (primeralph):

Give that guy a cookie.

OpenStudy (anonymous):

2 cookies!